OpenStudy (anonymous):

At your retail store, a customer is purchasing a new blood pressure monitor. The regular price is $100.00, but it's on sale this week for 25% 0ff. Additionally, today is senior Tuesday, and they get an additional 5% off. What will the final price be after discounts and 5% sales tax?

OpenStudy (anonymous):

$74.81 100.00* (.25) = 25.00 Subtract from regular price : 100.00 -25.00=75.00 75.00 * (.05) =3.75 Subtract from price : 75-3.75 = 71.25 71.25 * (1.05) = $74.81